# Chicago, IL 60607-7045 USA

### v2

@@ Abstract.
This is a two-part article. In the first part, we study an alternative notion to Nagata rings. A Nagata ring is a Noetherian ring $R$ such that every finite $R$-algebra that is an integral domain has finite normalization. We replace the normalization by an $(S_{2})$-ification, study new phenomena, and prove parallel results. In particular, we show a Nagata domain has a finite $(S_{2})$-ification. In the second part, we study the local lifting problem. We show that for a semilocal Noetherian ring $R$ that is $I$-adically complete for an ideal $I$, if $R/I$ has $(S_{k})$ (resp. Cohen–Macaulay, Gorenstein, lci) formal fibers, so does $R$. As a consequence, we show if $R/I$ is a quotient of a Cohen–Macaulay ring, so is $R$.
@@ 1. Introduction
In previous work (see [Lyu, §8]), the author showed that $\mathbf{R}$=“is a quotient of a Gorenstein ring” satisfies the lifting property. However, whether or not $\mathbf{R}$=“is a quotient of a Cohen–Macaulay ring” satisfies the lifting property seems to be difficult. We provide two perspectives on this question.
@@ 4. $(iv)$
Note that when $k=1$, this is exactly the semi-Nagata property (Theorem 1.2(i)(i)).
@@
In our case, such an assignment of ideals can be found with effort. We do a basic reduction in §10, saying we just need to assign $\mathfrak{m}$-primary ideals to complete local rings $(A,\mathfrak{m})$ that are $\mathbf{P}$ exactly on the punctured spectrum, strictly functorial with respect to flat maps with $\mathbf{P}$-fibers and 0-dimensional special fiber. After the reduction we find a desired assignment for all properties but lci in §11. The case $\mathbf{R}$=“is a quotient of a Cohen–Macaulay ring” of Theorem 1.8 follows from the case $\mathbf{R}$=“has Cohen–Macaulay formal fibers” via the argument already present in [Lyu, §8].
@@ 4. Naive and canonical $(S_{2})$-closures
Assume $R$ is semilocal. For $P\in\operatorname{OS}_{2}^{\wedge}({R})$, the punctured spectrum of the ring $(R^{\wedge})_{\mathfrak{P}}$ is disconnected, as it has an isolated point given by a minimal prime $P_{0}\subsetneq P$ with $\operatorname{ht}(P/P_{0})=1$. By [Stacks, Tag 0BLR] we have $0pt(R^{\wedge})_{\mathfrak{P}}<2$, so $0ptR_{P\cap R}<2$. Therefore, if $R$ is $(S_{2})$, then $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$.
@@ 4. $(iv)$
Observe that for every $\mathfrak{p}\in\operatorname{Spec}_{1}(R)$, we have $R_{\mathfrak{p}}=(R^{n\sigma})_{\mathfrak{p}}$. As $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$, we see that for every $P\in\operatorname{Spec}(R^{\wedge})$ so that $\operatorname{ht}(P/P_{j})=1$ for some $1\leq j\leq m$, we have $(R^{\wedge})_{\mathfrak{P}}=(R^{n\sigma}\otimes_{R}R^{\wedge})_{\mathfrak{P}}$. Therefore the sub-$S_{j}$-algebra $T_{j}$ of $L_{j}$ satisfies $(S_{j})_{P}=(T_{j})_{\mathfrak{P}}$ for all $P\in\operatorname{Spec}_{1}(S_{j})$.
@@
Let $A$ be a Noetherian complete local ring that is $(S_{1})$ and has irreducible spectrum. Let $F$ be the total fraction ring of $A$ and let $M\subseteq N$ be two sub-$A$-modules of $F$. If $M_{P}=N_{\mathfrak{P}}$ for all $P\in\operatorname{Spec}_{1}(A)$, and $M$ is finite, then $N$ is finite.
@@
Let $U$ be the locus where $M$ is $(S_{2})$. Then $U$ is open [EGA ̵IV2, Proposition 6.11.6], and contains $\operatorname{Spec}_{1}(A)$ as $A$ is $(S_{1})$. Let $j:U\to\operatorname{Spec}(A)$ be the canonical open immersion. It is clear that $j_{*}j^{*}M=\bigcap_{P\in\operatorname{Spec}_{1}(A)}M_{\mathfrak{P}}$.
Let $P_{0}$ be the minimal prime of $A$, and let $\overline{A}=A/P_{0}$. Note that $\operatorname{Ass}_{A}(M)=\{P_{0}\}$. For $P\in\operatorname{Spec}(A)\setminus U$, the completion $\overline{A}^{\wedge}_{\mathfrak{P}}$ is $(S_{1})$ as $\overline{A}$ is $(S_{1})$ with $(S_{1})$ formal fibers. Moreover, by Ratliff [Stacks, Tag 0AW6], $\overline{A}^{\wedge}_{\mathfrak{P}}$ is equidimensional, as $\overline{A}_{\mathfrak{P}}$ is equidimensional and universally catenary. Therefore for all $\mathfrak{p}\in\operatorname{Ass}(\overline{A}^{\wedge}_{\mathfrak{P}})$ we have $\dim\overline{A}^{\wedge}_{\mathfrak{P}}/\mathfrak{p}=\operatorname{ht}(P)>1$. By [Stacks, Tag 0BK3], $j_{*}j^{*}M$ is finite; and so is its submodule $N$. ∎
@@ 5. Extension rings via first local cohomology
By standard theory of limits, particularly [Stacks, Tags 01ZO and 081E], we may assume $W$ is generated by a single element $f$. Let $T$ be a finite $(W^{-1}U)$-modification of $W^{-1}R$. Then $T$ and $\mathcal{O}_{U}$ glue to a coherent $\mathcal{O}_{V}$-algebra $\mathcal{A}$, where $V=U\cup D(f)$. Let $j:V\to\operatorname{Spec}(A)$ be the open immersion, and let $\mathcal{A}^{\prime}$ be the integral closure of $\mathcal{O}_{\operatorname{Spec}(A)}$ in $j_{*}\mathcal{A}$. Then $j^{*}\mathcal{A}^{\prime}=\mathcal{A}$, and we can find a coherent sub-$\mathcal{O}_{\operatorname{Spec}(A)}$-algebra $\mathcal{A}^{\prime\prime}$ of $\mathcal{A}^{\prime}$ such that $j^{*}\mathcal{A}^{\prime\prime}=\mathcal{A}$, which corresponds to a $U$-modification $R^{\prime\prime}$ of $R$ that satisfies $W^{-1}R^{\prime\prime}=T$. Alternatively, one can use Zariski’s Main Theorem, [Stacks, Tag 05K0]. ∎
@@
Let $(R,\mathfrak{m})$ be a Noetherian local domain of depth at least $1$ and dimension at least $2$. Assume that $\mathfrak{m}\in\operatorname{Spec}(R)$ is p-unibranch.
@@
Let $M_{1}$ be the socle of $H^{1}_{\mathfrak{m}}(R)$, which is finite as $H^{1}_{\mathfrak{m}}(R)$ is Artinian [Har66, Chapter V, Corollary 6.5]. The ring $R_{1}:=R[\mathfrak{m};M_{1}]$ is then finite over $R$ by Lemma 5.5. As $\mathfrak{m}\in\operatorname{Spec}(R)$ is p-unibranch, $R_{1}$ is a local ring, and we denote its maximal ideal by $\mathfrak{m}_{1}$. We know $\dim R_{1}=\dim R\geq 2$, and we know $0ptR_{1}\geq 1$ as $R_{1}$ is a subring of $\mathcal{O}(D(\mathfrak{m}))$. Therefore the same applies to the ring $R_{1}$, and from Discussion 5.3 we see $R[\mathfrak{m};M_{2}]$ is finite over $R$ for $M_{2}=(M_{1}^{a}:\mathfrak{m})$. Inductively we see $R[\mathfrak{m};H^{1}_{\mathfrak{m}}(R)]$ is integral over $R$, and so is its subring $R[\mathfrak{m};M]$. ∎
@@ 6. Eliminating the obstructions
By Theorem 3.1, there exists a finite $D(\mathfrak{p})$-modification $R_{1}$ of $R$ such that all preimages of $\mathfrak{p}$ in $\operatorname{Spec}(R_{1})$ are p-unibranch. We know $R_{1}\subseteq R^{\sigma}$ as $\operatorname{ht}(\mathfrak{p})>1$. Next, we take a finite $D(\mathfrak{p}R_{1})$-modification $R_{2}$ of $R_{1}$, so that the number of maximal ideals of $(R_{2}^{\wedge})_{\mathfrak{P}}$ (i.e. the number of preimages of $P$ in $\operatorname{Spec}(R_{2}^{\wedge})$) is maximal among all possible $R_{2}$; to see this is achievable, note that $(R_{2}^{\wedge})_{\operatorname{red}}$ is finite birational over $(R_{1}^{\wedge})_{\operatorname{red}}$, so $((R_{2}^{\wedge})_{\operatorname{red}})_{\mathfrak{P}}$ is contained in the normalization of $((R_{1}^{\wedge})_{\operatorname{red}})_{\mathfrak{P}}$, which is finite as a complete Noetherian local ring is Nagata [Stacks, Tag 0335], so the number of maximal ideals of $(R_{2}^{\wedge})_{\mathfrak{P}}$ is bounded. Note that we still have $R_{2}\subseteq R^{\sigma}$ and that all preimages of $\mathfrak{p}$ in $\operatorname{Spec}(R_{2})$ are p-unibranch. We will show no FONSIs of $R_{2}$ are above $P$.
@@ 3. $(iii)$
On the other hand, if $R$ is a Noetherian universally catenary domain and $R^{\prime}$ is an integral domain containing and finite over $R$, then $R^{\prime}$ is $(S_{2})$ as a ring if and only if $R^{\prime}$ as an $R$-module. To see this, let us assume $(R,\mathfrak{m})$ local, so $R^{\prime}$ is semilocal with maximal ideals $\mathfrak{m}^{\prime}_{1},\ldots,\mathfrak{m}^{\prime}_{n}$. Since $R$ is universally catenary we know $\operatorname{ht}(\mathfrak{m}^{\prime}_{i})=\dim R$ for every $i$, see [Stacks, Tag 02IJ]. Moreover, as $\mathfrak{m}^{\prime}_{i}$ are the only preimages of $\mathfrak{m}$ in $\operatorname{Spec}(R^{\prime})$, we have $\min_{i}0ptR^{\prime}_{\mathfrak{m}^{\prime}_{i}}=0pt_{R}R^{\prime}$. We conclude that $0pt_{R}R^{\prime}\geq\max\{2,\dim R\}$ if and only if $0ptR^{\prime}_{\mathfrak{m}^{\prime}_{i}}\geq\max\{2,\operatorname{ht}(\mathfrak{m}^{\prime}_{i})\}$ for all $i$. This gives the original assertion via localization.
@@ 8. Lifting the semi-Nagata property
As $M$ is a submodule of $K$, we know $M\subseteq M_{\mathfrak{p}}$ for all $\mathfrak{p}\in\mathcal{M}$, in particular $M$ is $f$-adically separated. Therefore $M$ is a submodule of its $f$-adic completion, which is finite as $R$ is $f$-adically complete and $M/fM$ is finite.
@@ 9. The local lifting argument
A Noetherian ring $A$ is said to be a $\mathbf{P}$-ring if its formal fibers are geometrically $\mathbf{P}$, in other words, $A_{\mathfrak{p}}\to A^{\wedge}_{\mathfrak{p}}$ is a $\mathbf{P}$-map for all $\mathfrak{p}\in\operatorname{Spec}(A)$. By [Stacks, Tag 0BIU], if $\mathbf{P}$ satisfies (I)(III)(IV) below, then a Noetherian ring $A$ is a $\mathbf{P}$-ring if and only if $A_{\mathfrak{m}}\to A^{\wedge}_{\mathfrak{m}}$ is a $\mathbf{P}$-map for all $\mathfrak{m}\in\operatorname{Max}(A)$. When $A$ is semilocal, this is to say $A\to A^{\wedge}$ is a $\mathbf{P}$-map (cf. [EGA ̵IV2, Proposition 7.3.14]); and if $\mathbf{P}$ satisfies (I)(II)(III)(IV) below, this is also equivalent to that for every finite $A$-algebra $B$ that is an integral domain, $(B^{\circ})^{-1}B^{\wedge}$ satisfies $\mathbf{P}$, see [EGA ̵IV2, Proposition 7.3.16]. By [Stacks, Tag 0BIV], an essentially finitely generated algebra over a $\mathbf{P}$-ring is a $\mathbf{P}$-ring.
@@
Whether or not $\mathbf{P}$ satisfies (VI) is generally called the Grothendieck localization problem for $\mathbf{P}$. The paper [Mur22] provides a uniform treatment to this problem, and provides a list of references of known results on what properties satisfy (I)(VI). In particular, (I)(VI) hold for $\mathbf{P}$$(S_{k})$,” “Cohen–Macaulay,” “Gorenstein,” and “lci.”
@@
We know $(R^{\wedge})_{\mathfrak{p}}$ is (quasi-)excellent as $R^{\wedge}$ is complete, therefore $((R^{\wedge})_{\mathfrak{p}})^{*}$ is quasi-excellent [KS21]. In particular, both $(R^{\wedge})_{\mathfrak{p}}$ and $((R^{\wedge})_{\mathfrak{p}})^{*}$ are $\mathbf{P}$-rings. By (8) $(R_{\mathfrak{p}})^{*}$ is also.
@@
If lci implies $\mathbf{P}$, then every lci ring is a $\mathbf{P}$-ring (cf. [Avr99, (5.4)]). The ring $((R^{\wedge})_{\mathfrak{p}})^{*}$ is a quotient of a regular ring as $R^{\wedge}$ is, so it is a $\mathbf{P}$-ring, avoiding [KS21]. This is the case in our applications.

### v3

@@ Abstract.
This is a two-part article. In the first part, we study an alternative notion to Nagata rings. A Nagata ring is a Noetherian ring $R$ such that every finite $R$-algebra that is an integral domain has finite normalization. We replace the normalization by an $(S_{2})$-ification, study new phenomena, and prove parallel results. In particular, we show a Nagata domain has a finite $(S_{2})$-ification. In the second part, we study the local lifting problem. We show that for a semilocal Noetherian ring $R$ that is $I$-adically complete for an ideal $I$, if $R/I$ has $(S_{k})$ (resp. Cohen–Macaulay, Gorenstein, lci) formal fibers, so does $R$. As a consequence, we show if $R/I$ is a quotient of a Cohen–Macaulay ring, so is $R$. We also discuss difficulties in lifting geometrically $(R_{k})$ formal fibers.
@@ 1. Introduction
In previous work (see [Lyu25a, §8]), the author showed that $\mathbf{R}$=“is a quotient of a Gorenstein ring” satisfies the lifting property. However, whether or not $\mathbf{R}$=“is a quotient of a Cohen–Macaulay ring” satisfies the lifting property seems to be difficult. We provide two perspectives on this question.
@@ 4. $(iv)$
Note that when $k=1$, having $(S_{k})$ formal fibers is exactly the semi-Nagata property (Theorem 1.2(i)(i)).
@@
In our case, such an assignment of ideals can be found with effort. We do a basic reduction in §10, saying we just need to assign $\mathfrak{m}$-primary ideals to complete local rings $(A,\mathfrak{m})$ that are $\mathbf{P}$ exactly on the punctured spectrum, strictly functorial with respect to flat maps with $\mathbf{P}$-fibers and 0-dimensional special fiber. After the reduction we find a desired assignment for all properties but lci in §11. The case $\mathbf{R}$=“is a quotient of a Cohen–Macaulay ring” of Theorem 1.8 follows from the case $\mathbf{R}$=“has Cohen–Macaulay formal fibers” via the argument already present in [Lyu25a, §8].
@@
Finally, in §14, we treat the lifting of $\mathbf{P}$ in codimension 0” and “Cohen–Macaulay in codimension 11.” This section mainly serves as a discussion of difficulties running the arguments in §§913 for other standard properties such as $(R_{k})$, but some positive results are obtained.
@@ 4. Naive and canonical $(S_{2})$-closures
Assume $R$ is semilocal. For $P\in\operatorname{OS}_{2}^{\wedge}({R})$, the punctured spectrum of the ring $(R^{\wedge})_{\mathfrak{P}}$ is disconnected, as it has an isolated point given by a minimal prime $P_{0}\subsetneq P$ with $\operatorname{ht}(P/P_{0})=1$. By [Stacks, Tag 0BLR] we have $0pt(R^{\wedge})_{P}<2$, so $0ptR_{P\cap R}<2$. Therefore, if $R$ is $(S_{2})$, then $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$.
@@ 4. $(iv)$
Observe that for every $\mathfrak{p}\in\operatorname{Spec}_{1}(R)$, we have $R_{\mathfrak{p}}=(R^{n\sigma})_{\mathfrak{p}}$. As $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$, we see that for every $P\in\operatorname{Spec}(R^{\wedge})$ so that $\operatorname{ht}(P/P_{j})=1$ for some $1\leq j\leq m$, we have $(R^{\wedge})_{P}=(R^{n\sigma}\otimes_{R}R^{\wedge})_{P}$. Therefore the sub-$S_{j}$-algebra $T_{j}$ of $L_{j}$ satisfies $(S_{j})_{P}=(T_{j})_{P}$ for all $P\in\operatorname{Spec}_{1}(S_{j})$.
@@
Let $A$ be a Noetherian complete local ring that is $(S_{1})$ and has irreducible spectrum. Let $F$ be the total fraction ring of $A$ and let $M\subseteq N$ be two sub-$A$-modules of $F$. If $M_{P}=N_{P}$ for all $P\in\operatorname{Spec}_{1}(A)$, and $M$ is finite, then $N$ is finite.
@@
Let $U$ be the locus where $M$ is $(S_{2})$. Then $U$ is open [EGA ̵IV2, Proposition 6.11.6], and contains $\operatorname{Spec}_{1}(A)$ as $A$ is $(S_{1})$. Let $j:U\to\operatorname{Spec}(A)$ be the canonical open immersion. It is clear that $j_{*}j^{*}M=\bigcap_{P\in\operatorname{Spec}_{1}(A)}M_{P}$.
Let $P_{0}$ be the minimal prime of $A$, and let $\overline{A}=A/P_{0}$. Note that $\operatorname{Ass}_{A}(M)=\{P_{0}\}$. For $P\in\operatorname{Spec}(A)\setminus U$, the completion $\overline{A}^{\wedge}_{P}$ is $(S_{1})$ as $\overline{A}$ is $(S_{1})$ with $(S_{1})$ formal fibers. Moreover, by Ratliff [Stacks, Tag 0AW6], $\overline{A}^{\wedge}_{P}$ is equidimensional, as $\overline{A}_{P}$ is equidimensional and universally catenary. Therefore for all $\mathfrak{p}\in\operatorname{Ass}(\overline{A}^{\wedge}_{P})$ we have $\dim\overline{A}^{\wedge}_{P}/\mathfrak{p}=\operatorname{ht}(P)>1$. By [Stacks, Tag 0BK3], $j_{*}j^{*}M$ is finite; and so is its submodule $N$. ∎
@@ 5. Extension rings via first local cohomology
By standard theory of limits, particularly [Stacks, Tags 01ZO and 081E], we may assume $W$ is generated by a single element $f$. Let $T$ be a finite $(W^{-1}U)$-modification of $W^{-1}R$. Then $T$ and $\mathcal{O}_{U}$ glue to a coherent $\mathcal{O}_{V}$-algebra $\mathcal{A}$, where $V=U\cup D(f)$. Let $j:V\to\operatorname{Spec}(A)$ be the open immersion, and let $\mathcal{A}^{\prime}$ be the integral closure of $\mathcal{O}_{\operatorname{Spec}(A)}$ in $j_{*}\mathcal{A}$. Then $j^{*}\mathcal{A}^{\prime}=\mathcal{A}$, and we can find a coherent subalgebra $\mathcal{A}^{\prime\prime}$ of $\mathcal{A}^{\prime}$ such that $j^{*}\mathcal{A}^{\prime\prime}=\mathcal{A}$, which corresponds to a $U$-modification $R^{\prime\prime}$ of $R$ that satisfies $W^{-1}R^{\prime\prime}=T$. Alternatively, one can use Zariski’s Main Theorem, [Stacks, Tag 05K0]. ∎
@@
Let $(R,\mathfrak{m})$ be a Noetherian local domain of dimension at least $2$. Assume that $\mathfrak{m}\in\operatorname{Spec}(R)$ is p-unibranch.
@@
Let $M_{1}$ be the socle of $H^{1}_{\mathfrak{m}}(R)$, which is finite as $H^{1}_{\mathfrak{m}}(R)$ is Artinian [Har66, Chapter V, Corollary 6.5]. The ring $R_{1}:=R[\mathfrak{m};M_{1}]$ is then finite over $R$ by Lemma 5.5. As $\mathfrak{m}\in\operatorname{Spec}(R)$ is p-unibranch, $R_{1}$ is a local ring, and we denote its maximal ideal by $\mathfrak{m}_{1}$. We know $\dim R_{1}=\dim R\geq 2$, therefore the same argument applies to the ring $R_{1}$, and from Discussion 5.3 we see $R[\mathfrak{m};M_{2}]$ is finite over $R$ for $M_{2}=(M_{1}^{a}:\mathfrak{m})$. Inductively we see $R[\mathfrak{m};H^{1}_{\mathfrak{m}}(R)]$ is integral over $R$, and so is its subring $R[\mathfrak{m};M]$. ∎
@@ 6. Eliminating the obstructions
By Theorem 3.1, there exists a finite $D(\mathfrak{p})$-modification $R_{1}$ of $R$ such that all preimages of $\mathfrak{p}$ in $\operatorname{Spec}(R_{1})$ are p-unibranch. We know $R_{1}\subseteq R^{\sigma}$ as $\operatorname{ht}(\mathfrak{p})>1$. Next, we take a finite $D(\mathfrak{p}R_{1})$-modification $R_{2}$ of $R_{1}$, so that the number of maximal ideals of $(R_{2}^{\wedge})_{P}$ (i.e. the number of preimages of $P$ in $\operatorname{Spec}(R_{2}^{\wedge})$) is maximal among all possible $R_{2}$; to see this is achievable, note that $(R_{2}^{\wedge})_{\operatorname{red}}$ is finite birational over $(R_{1}^{\wedge})_{\operatorname{red}}$, so $((R_{2}^{\wedge})_{\operatorname{red}})_{P}$ is contained in the normalization of $((R_{1}^{\wedge})_{\operatorname{red}})_{P}$, which is finite as a complete Noetherian local ring is Nagata [Stacks, Tag 0335], so the number of maximal ideals of $(R_{2}^{\wedge})_{P}$ is bounded. Note that we still have $R_{2}\subseteq R^{\sigma}$ and that all preimages of $\mathfrak{p}$ in $\operatorname{Spec}(R_{2})$ are p-unibranch. We will show no FONSIs of $R_{2}$ are above $P$.
@@ 3. $(iii)$
On the other hand, if $R$ is a Noetherian universally catenary domain and $R^{\prime}$ is an integral domain containing and finite over $R$, then $R^{\prime}$ is $(S_{2})$ as a ring if and only if $R^{\prime}$ is $(S_{2})$ as an $R$-module. To see this, let us assume $(R,\mathfrak{m})$ local, so $R^{\prime}$ is semilocal with maximal ideals $\mathfrak{m}^{\prime}_{1},\ldots,\mathfrak{m}^{\prime}_{n}$. Since $R$ is universally catenary we know $\operatorname{ht}(\mathfrak{m}^{\prime}_{i})=\dim R$ for every $i$, see [Stacks, Tag 02IJ]. Moreover, as $\mathfrak{m}^{\prime}_{i}$ are the only preimages of $\mathfrak{m}$ in $\operatorname{Spec}(R^{\prime})$, we have $\min_{i}0ptR^{\prime}_{\mathfrak{m}^{\prime}_{i}}=0pt_{R}R^{\prime}$. We conclude that $0pt_{R}R^{\prime}\geq\max\{2,\dim R\}$ if and only if $0ptR^{\prime}_{\mathfrak{m}^{\prime}_{i}}\geq\max\{2,\operatorname{ht}(\mathfrak{m}^{\prime}_{i})\}$ for all $i$. This gives the original assertion via localization.
@@ 8. Lifting the semi-Nagata property
As $M$ is a submodule of $K$, we know $M\subseteq M_{\mathfrak{p}}=R_{\mathfrak{p}}$ for all $\mathfrak{p}\in\mathcal{M}$, in particular $M$ is $f$-adically separated. Therefore $M$ is a submodule of its $f$-adic completion, which is finite as $R$ is $f$-adically complete and $M/fM$ is finite.
@@ 9. The local lifting argument
A Noetherian ring $A$ is said to be a $\mathbf{P}$-ring if its formal fibers are geometrically $\mathbf{P}$, in other words, $A_{\mathfrak{p}}\to A^{\wedge}_{\mathfrak{p}}$ is a $\mathbf{P}$-map for all $\mathfrak{p}\in\operatorname{Spec}(A)$. By [Stacks, Tag 0BIU], if $\mathbf{P}$ satisfies (I)(IV) below, then a Noetherian ring $A$ is a $\mathbf{P}$-ring if and only if $A_{\mathfrak{m}}\to A^{\wedge}_{\mathfrak{m}}$ is a $\mathbf{P}$-map for all $\mathfrak{m}\in\operatorname{Max}(A)$. When $A$ is semilocal, this is to say $A\to A^{\wedge}$ is a $\mathbf{P}$-map (cf. [EGA ̵IV2, Proposition 7.3.14]); and if $\mathbf{P}$ satisfies (I)(IV) below, this is also equivalent to that for every finite $A$-algebra $B$ that is an integral domain, $(B^{\circ})^{-1}B^{\wedge}$ satisfies $\mathbf{P}$, see [EGA ̵IV2, Proposition 7.3.16]. By [Stacks, Tag 0BIV], an essentially finitely generated algebra over a $\mathbf{P}$-ring is a $\mathbf{P}$-ring.
@@
Whether or not $\mathbf{P}$ satisfies (VI) is generally called the Grothendieck localization problem for $\mathbf{P}$. The paper [Mur22] provides a uniform treatment of this problem, and provides a list of references of known results on what properties satisfy (I)(VI). In particular, (I)(VI) hold for $\mathbf{P}$$(S_{k})$,” “Cohen–Macaulay,” “Gorenstein,” and “lci.”
@@
We know $(R^{\wedge})_{\mathfrak{p}}$ is (quasi-)excellent as $R^{\wedge}$ is complete, therefore $((R^{\wedge})_{\mathfrak{p}})^{*}$ is quasi-excellent [KS21]. In particular, both $(R^{\wedge})_{\mathfrak{p}}$ and $((R^{\wedge})_{\mathfrak{p}})^{*}$ are $\mathbf{P}$-rings and have open $\mathbf{P}$-locus. By (8) $(R_{\mathfrak{p}})^{*}$ is a local $\mathbf{P}$-ring, hence its $\mathbf{P}$-locus is open by (V), cf. [Stacks, Tag 02JY].
@@
If lci implies $\mathbf{P}$, then every lci ring is a $\mathbf{P}$-ring (cf. [Avr99, (5.4)]). The ring $((R^{\wedge})_{\mathfrak{p}})^{*}$ is a quotient of a regular ring as $R^{\wedge}$ is, so it is a $\mathbf{P}$-ring, avoiding [KS21]. This is the case in our applications in §13.